lenght of the shadow of a 15m high pole is 5sqrt3 m at 7o’ clock in the morning . then what is the angle of elevation of sun rays with the ground at the time 

Grade:12th pass

2 Answers

Arun
5 years ago
Let us take the scenario in the form of a right triangle, where the height of the pole is represented as AB, the shadow of the pole is represented as BC. Let 
∠ C be the angle of Elevation.
So, We know that,
Tan Ф = Opposite / Adjacent
=> Tan Ф = 15 / 15 √ 3
=> Tan Ф = 1 / √ 3
We know that, Tan 30° = 1 / √ 3
So we can say that, Ф = 30°
Hence the angle of Elevation of the sum at that time was 30°.
Sripad Sambrani
4 years ago
Given:
height of pole =15m
length of shadow=5√3
 
RTF: angle of sun’s elevation
 
Solution:
Let θ be the angle of sun’s elevation
Tanθ=15/5√3=√3=tan(π/3)
Hence angle of sun’s elevation is π/3